the cartesian coordinates of a point are given by (x,y)=(-2.5,-3.5) meters. Find the polar coordinates of this point?


1
Expert's answer
2021-07-05T08:41:33-0400

The polar coordinates are given as follows:


"r = \sqrt{x^2 + y^2}\\
\theta = \arctan\dfrac{y}{x}"

where "x,y" are the cartesian coordinates. Thus, obtain:

"r = \sqrt{(-2.5)^2 + (-3.5)^2} \approx 4.3\\
\theta = \arctan\dfrac{-3.5}{-2.5} \approx 234\degree"